# iclr2025_conference.bib

### v3

@@ Abstract
As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization of training data has attracted growing attention. Existing works in this direction aim to establish an understanding of whether or to what extent DPMs learn via memorization. Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI. Existing works revealed that conditional DPMs are more prone to training data memorization than unconditional DPMs, and the motivated data extraction methods are mostly for conditional DPMs. However, these understandings are primarily empirical, and extracting training data from unconditional models has been found to be extremely challenging. In this work, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs under the assumption of model convergence. Our theoretical analysis indicates that extracting data from unconditional models can also be effective by constructing a proper surrogate condition. Based on this result, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on the generated data as a surrogate condition to extract training data from unconditional DPMs. Empirical results demonstrate that our SIDE can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset.
@@ 1 Introduction
The diffusion probabilistic models (DPMs) ho2020denoising; sohl2015deep; song2019generative is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, and Sora videoworldsimulators2024. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, biased generation, privacy breaches, or copyright infringement, as highlighted in the literature Somepalli2022DiffusionAO; somepalli2023understanding; asay2020independent; cooper2024files; perera2023analyzing. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training databutterick2023stable. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by carlini2023extracting demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion rombach2022high, revealing the potential dangers associated with these models.
@@
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., being conditional is more prone to memorization somepalli2023understanding. Gu2023OnMI investigates the influential factors on memorization behaviors via a comprehensive set of experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though inspiring, these understandings are primarily empirical. Moreover, without a unified theoretical understanding of memorization in both conditional and unconditional DPMs, it is extremely difficult to design an effective data extraction method for unconditional DPMs.
@@
In this work, we first introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner. Based on this metric, we present a theoretical framework that explains why conditional generative models memorize more data and why random labeling can lead to increased memorization. Our theoretical analysis indicates that a classifier trained on the same or similar training data can serve as a surrogate condition for unconditional DPMs. By further making the classifier time-dependent (to the diffusion sampling process), we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs. We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1), confirming the accuracy of the theoretical framework.
@@
  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical framework that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) implicit labels (e.g., the learned clusters) can serve as a surrogate condition for unconditional DPMs.
  • Based on our theoretical understanding, we propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets, and show that, on average, it can outperform the baseline method proposed by carlini2023extracting by more than 50%.
@@ 2 Related Work
Diffusion Probabilistic Models  DPMs sohl2015deep such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, Sora videoworldsimulators2024, Runway rombach2022high, and Imagen saharia2022photorealistic have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks dhariwal2021diffusion. These models can be viewed from two perspectives. The first is score matching song2019generative, where diffusion models learn the gradient of the image distribution song2020score. The second perspective involves denoising DPMs ho2020denoising, which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, dhariwal2021diffusion utilizes a classifier to guide the denoising process at each sampling step. Additionally, ho2022classifier introduces classifier-free guidance for conditional data sampling using DPMs.
@@
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models carlini2022quantifying; jagielski2022measuring, which has motivated more in-depth research on DPMs somepalli2023understanding; Gu2023OnMI; rahman2024frame. Notably, a recent research on image-generation DPMs: Somepalli2022DiffusionAO found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by carlini2023extracting in broader experiments for both conditional and unconditional diffusion models. Further studies somepalli2023understanding and Gu2023OnMI linked memorization to model conditioning, showing that conditional models are more prone to memorization. Several works have proposed methods to detect and mitigate memorization to address this issue. Wen et al. wen2024detecting proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions. It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization. Ren et al. ren2024unveiling proposed two metrics based on findings on cross-attention in DPMs to detect memorization. To mitigate memorization, Chen et al.chen2024towards proposed anti-memorization guidance to mitigate memorization during the sampling process. Instead of using guidance, Ren et al. ren2024unveiling mitigate memorization by changing the attention score or masking out summary tokens in the cross-attention layer. Wen et al.wen2024detecting do so by minimizing the magnitude during inference. So far, the motivated data extraction attack or defend methods from these empirical understandings are mostly for conditional DPMs carlini2023extracting; webster2023reproducible, and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs Gu2023OnMI. Although the empirical understandings are inspiring, a theoretical understanding of the memorization behaviors of DPMs is missing from the current literature. A recent attempt by ross2024a uses the local intrinsic dimension (LID) metric to characterize memorization in DPMs. However, the finding that lower LID leads to more memorization is only verified on a toy 1-dimensional dataset and only a few generated images, and it fails to address the data extraction challenge on unconditional DPMs. This challenge is exemplified by the brute-force methods currently employed, as discussed in Somepalli2022DiffusionAO; carlini2023extracting, highlighting the inadequacies of existing approaches for extracting data from unconditional diffusion models. In response to these challenges, we introduce a theoretical framework to characterize memorization in DPMs, which further motivates a novel data extraction method for unconditional DPMs.
3 Proposed Theory
In this section, we introduce a memorization metric and provide a theoretical explanation for the universality of data memorization in both conditional and unconditional diffusion models.
3.1 Memorization Metric
Intuitively, the memorization of fixed training data points (i.e., pointwise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point. Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
Definition 1 (Pointwise Memorization)
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
${\mathcal{M}}_{point}({\mathcal{D}};\theta)=\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}},$ (1)
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ is the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make equation (1) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could use the Gaussian distribution with a covariance matrix $\epsilon\textit{I}$ (I is the identity matrix) for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$:
$q({\bm{x}},{\bm{x}}_{i},\epsilon)=\frac{1}{\sqrt{(2\pi\epsilon)^{d}}}\exp\left\{-\frac{1}{2\epsilon}({\bm{x}}-{\bm{x}}_{i})^{\top}({\bm{x}}-{\bm{x}}_{i})\right\}.$ (2)
Note that in equation (1), a smaller value of ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the intrinsic memorization capability of model $p_{\theta}$.
Semantic Memorization
As a metric, ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ should be monotonous with the actual semantic memorization effect, which refers to the model’s tendency to reproduce unique features of the training samples rather than generating near-duplicate examples. Here, we define semantic memorization in the latent space as it distills the unique semantic information of each training sample.
Definition 2 (Semantic Memorization)
Let ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ be the dataset, $p_{\theta}(x)$ be the PDF of the generated samples by model $p_{\theta}$, and $\bm{z}$ be the learned latent code for data sample $\bm{x}$. The semantic memorization of model $p_{\theta}$ on ${\mathcal{D}}$ is:
[…]
where $\bm{z_{i}}$ denotes the ground truth latent code of $\bm{x_{i}}$.
${\mathcal{M}}_{semantic}\left({\mathcal{D}};\theta\right)$ measures semantic memorization because it evaluates how well the learned representations ${\bm{z}}$ of training samples align with their ground truth latent codes, capturing the underlying structure of the data rather than simply memorizing the exact duplicates of the training data points. In practice, the ground truth latent codes are known but can be approximated by an independent encoder.
The following theorem formulates the relationship between pointwise and semantic memorization.
Theorem 1
Pointwise memorization ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ is monotonic to semantic memorization, formally:
$\frac{\partial{\mathcal{M}}_{point}({\mathcal{D}};\theta)}{\partial{\mathcal{M}}_{semantic}({\mathcal{D}};\theta)}=\frac{1}{2\epsilon}+\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2}>0$ (4)
where $\Sigma_{p_{\theta}}$ is the covariance matrix of the learned latent distribution of the training data.
We first build the relationship map ${\mathcal{M}}_{point}$ into latent space, and then break down equation (4) into separate components and derive each component separately. So, the equation (4) can be simplified to $\frac{1}{2\epsilon}$. The detailed proof can be found in Appendix A.2. Note that the memorization effect studied in this work refers specifically to the pointwise memorization ${\mathcal{M}}_{point}$.
3.2 Theoretical Framework
Based on pointwise memorization, here we present a theoretically framework that explains why conditional DPMs memorize more data. Our theoretical framework is based on the concept of informative labels, which refers to information that can differentiate subsets of data samples. We first give a formal definition of informative labels and then prove their two key properties: 1) they facilitate tighter clustering of samples around their respective means, and 2) they reduce variance in the latent representations. Building upon the two properties, we theoretically show that conditional DPMs memorize more data.
Informative Labels
The concept of informative labels has been previously discussed as class labels Gu2023OnMI. In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We define an informative label as follows.
Definition 3 (Informative Label)
Let ${\mathcal{Y}}=\{y_{i},y_{2},\cdots,y_{C}\}$ be the label set for training dataset ${\mathcal{D}}$ with $C$ unique labels. $y_{i}$ is the associated label with ${\bm{x}}_{i}$, and ${\mathcal{D}}_{y=c}=\{{\bm{x}}_{i}:{\bm{x}}_{i}\in\mathcal{D},y_{i}=c\}$ is the subset of training samples shared the same label $y=c$. A label $y=c$ is an informative label if
$|{\mathcal{D}}_{y=c}|<|{\mathcal{D}}|.$ (5)
Here, the labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like silent features or clusters. Next, we will explore the correlation between informative labels and the clustering effect in the latent space of a generative model with an encoder and decoder. In diffusion models, the encoder represents the forward diffusion process, while the decoder represents the reverse process. Notably, the latent space can be defined at earlier timesteps according to ho2020denoising.
Memorization in Conditional DPMs
Informative labels cause a clustering effect in the latent space of DPMs by providing contextual information that allows the encoder to better differentiate between data samples. When data points are associated with informative labels, the encoder can map them to a latent distribution that accurately reflects their shared characteristics. This results in tighter clusters in the latent space as samples with the same informative label become more concentrated around their respective means. Consequently, the latent representations of these samples exhibit reduced variance, leading to a more structured and organized latent space. We prove these two properties under the assumption of model convergence.
Suppose we have an encoder $f_{\theta_{E}}({\bm{x}})$ and a decoder $f_{\theta_{D}}({\bm{z}})$. The encoder $f_{\theta_{E}}({\bm{x}})$ maps data samples ${\bm{x}}\in{\mathcal{D}}$ to the latent distribution ${\bm{z}}$, which follows a normal distribution ${\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$: $p_{\theta}({\bm{z}})={\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$. For ${\bm{x}}_{i}\in{\mathcal{D}}_{y=c}$, the encoder maps ${\bm{x}}_{i}$ to a latent distribution ${\bm{z}}_{c}$ subject to ${\mathcal{N}}(\mathbf{{\bm{\mu}}_{c}},\mathbf{\Sigma_{c}})$. The decoder $f_{\theta_{D}}({\bm{z}})$ maps ${\bm{z}}$ back to the original data samples ${\bm{x}}$. $y_{i}$ is the label of data sample ${\bm{x}}_{i}$. Training a generative $p_{\theta}$ is to optimize the following:
$\min_{\theta}-\sum_{{{\bm{x}}}_{\mathrm{i}}\in\mathrm{{\mathcal{D}}}}{\log{p}_{\theta}\left({{\bm{x}}_{i}}|y_{i}\right)}.$ (6)
Assumption 1
Given suffix training on ${\mathcal{D}}$, the generative model $p_{\theta}$ converges to an optimal solution for equation (6) : $\theta^{*}=\arg\min_{\theta}-\sum_{{\bm{x}}_{\mathrm{i}}\in\mathcal{D}}\log p_{\theta}({\bm{x}}_{i}|y_{i})$.
Based on the above assumption, we can derive the following Proposition 1, with detailed proof deferred to Appendix A.3
Proposition 1
Let ${\bm{z}}$ be the latent space of generative model $p_{\theta}$ conditioned on an informative label $y=c$, the latent representations learned by $p_{\theta}$ under Assumption 1 satisfy:
$\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})\leq\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}),$ (7)
$\|\mathbf{\Sigma}_{c}\|_{*}\leq\|\mathbf{\Sigma}\|_{*},$ (8)
where $\Sigma$ is the covariance matrix, $\Sigma_{c}$ is the covariance matrix conditioned on informative label $y=c$, $\bm{\mu}_{c}$ denotes the mean of the latent representations ${\bm{z}}$ conditioned on $y=c$ and $\bm{\mu}$ denotes the mean on the overall dataset ${\mathcal{D}}$.
Proposition 1 describes two properties of the learned latent space driven by informative labels: 1) tighter clustering as defined in equation (7) and 2) reduced variance as defined in equation (8). Tighter clustering means that the data samples associated with the same informative label are more closely clustered, which allows the model to more effectively capture and memorize the specific features and patterns relevant to those labels. This proximity in the latent space enhances the model’s ability to recall memorized samples during generation, as the representations are organized around distinct means. Additionally, reduced variance in these clustered representations leads to greater stability, ensuring that the model can consistently reproduce memorized outputs from the clustered latent codes. Based on this understanding, we formalize the relationship between informative labels and memorization in conditional DPMs via the following theorem.
Theorem 2
A generative model $p_{\theta}$ incurs a higher degree of pointwise memorization when conditioned on informative labels $y=c$, mathematically expressed as:
$\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta_{y=c})}{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta)}\leq 1$ (9)
where $\theta_{y=c}$ denotes the parameters of the model when trained on dataset ${\mathcal{D}}_{y=c}$.
The proof can be found in Appendix A.4. Theorem 2 states that any form of information labels can incur more memorization in DPMs, including conventional class labels and random labels. As shown in our proof, the informative nature of a label reduces the entropy (or uncertainty) of the data distribution conditioned on that label, leading to a more focused and memorable data representation. This explains the two empirical observations made in existing works Gu2023OnMI: 1) conditional DPMs are more prone to memorization and 2) even random labels can lead to more memorization. It also explains the findings that unconditional models do not replicate data and that text conditioning increases memorization Somepalli2022DiffusionAO; chen2024towards.
Memorization in Unconditional DPMs
According to Theorem 2, one could leverage informative labels to extract training data from conditional DPMs. Intuitively, the Text captions and class labels commonly used to train conditional DPMs are valid informative labels, which we call explicit informative labels. However, unconditional DPMs do not have explicit informative labels and thus are more difficult to extract training data from. Nevertheless, our theory indicates that the learned representation clusters by an unconditional DPM can also serve as a type of informative labels, which we call implict information labels. It means that if we can formulate the clustering information in the training data, we could construct implicit informative labels to help extract training data from unconditional DPMs. This motivates us to propose a new data extraction method SIDE for unconditional DPMs in the next section.

### v5

# appAppendix References

@@ Abstract
As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization has attracted growing attention. Existing works in this field aim to establish an understanding of whether or to what extent DPMs learn via memorization. Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI. Existing works revealed that conditional DPMs are more prone to memorize training data than unconditional DPMs. And most data extraction methods developed so far target conditional DPMs. Although unconditional DPMs are less prone to data extraction, further investigation into these attacks remains essential since they serve as the foundation for conditional models like Stable Diffusion, and exploring these attacks will enhance our understanding of memorization in DPMs. In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on generated data as surrogate conditions to extract training data from unconditional DPMs. Empirical results demonstrate that it can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset. Furthermore, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs and why SIDE is effective.
@@ 1 Introduction
The diffusion probabilistic models (DPMs) [16, 30, 34] is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion [26], DALL-E 3 [2], and Sora [3]. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches or copyright infringement, as highlighted in the literature [31, 32, 1, 11]. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data[4]. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by [5] demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion [26], revealing the potential privacy and copyright dangers associated with them.
@@
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., conditional-trained DPMs are more prone to memorization [32]. Gu et al. [13] investigates influential factors on memorization behaviors via comprehensive experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though these works indicate that unconditional diffusion models are much less prone to data extraction attacks, investigating such attacks for unconditional diffusion models is still necessary. This is because 1) unconditional DPMs are the foundations of conditional generative models like Stable Diffusion, which adapts pre-trained unconditional DPMs as the essential part for conditional generation or combines them with conditional training approaches [26, 39]; 2) exploring data extraction attacks on unconditional DPMs enhances understanding of memorization mechanisms in both conditional and unconditional DPMs.
@@
In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs. This attack method uses a classifier trained on the same or similar training data as a surrogate condition for unconditional DPMs. SIDE can extract training data effectively by making the classifier time-dependent (to the diffusion sampling process). We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1). Additionally, to offer a theoretical analysis, we introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner. Based on this metric, we explain why conditional DPMs are more likely to memorize training data, why random labelling can increase this memorization and why the SIDE method is effective for extracting training data in unconditional DPMs.
@@
  • We propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical analysis that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) why the SIDE method is effective for extracting training data in unconditional DPMs.
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets and show that, on average, it can outperform the baseline method by more than 50%.
@@ 2 Related Work
Diffusion Probabilistic Models  DPMs [30] such as Stable Diffusion [26], DALL-E 3 [2], Sora [3], Runway [26], and Imagen [28] have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks [12]. These models can be viewed from two perspectives. The first is score matching [34], where diffusion models learn the gradient of the image distribution [35]. The second perspective involves denoising DPMs [16], which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, [12] utilizes a classifier to guide the denoising process at each sampling step. Additionally, [17] introduces classifier-free guidance for conditional data sampling using DPMs.
@@
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models [6, 19], which has motivated more in-depth research on DPMs [32, 13, 24]. Notably, a recent research on image-generation DPMs: [31] found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by [5] in broader experiments for both conditional and unconditional diffusion models. Further studies [32] and [13] linked memorization to model conditioning, showing that conditional models are more prone to memorization. Several works have proposed methods to detect and mitigate memorization to address this issue. Wen et al. [38] proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions. It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization. Ren et al. [25] proposed two metrics based on findings on cross-attention in DPMs to detect memorization. To mitigate memorization, Chen et al.[7] proposed anti-memorization guidance to mitigate memorization during the sampling process. Instead of using guidance, Ren et al. [25] mitigated memorization by changing the attention score or masking out summary tokens in the cross-attention layer. Wen et al.[38] achieved this by minimizing the magnitude during inference. So far, the motivated data extraction attack or defend methods are primarily for conditional DPMs [5, 37], and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs [13]. Though challenging, investigating data extraction methods in unconditional DPMs is still necessary. This is because 1) Unconditional DPMs are key parts in conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) Exploring this approach enhances our understanding of memorization in DPMs. In response to these challenges, we propose a novel data extraction method called SIDE for unconditional DPMs and providing theoretical understanding of SIDE.